# AutoMisty: A Multi-Agent LLM Framework for Automated Code Generation in the Misty Social Robot

### v1

@@ Abstract
The social robot’s open API allows users to customize open-domain interactions. However, it remains inaccessible to those without programming experience. In this work, we introduce AutoMisty, the first multi-agent collaboration framework powered by large language models (LLMs), to enable the seamless generation of executable Misty robot code from natural language instructions. AutoMisty incorporates four specialized agent modules to manage task decomposition, assignment, problem-solving, and result synthesis. Each agent incorporates a two-layer optimization mechanism, with self-reflection for iterative refinement and human-in-the-loop for better alignment with user preferences. AutoMisty ensures a transparent reasoning process, allowing users to iteratively refine tasks through natural language feedback for precise execution. To evaluate AutoMisty’s effectiveness, we designed a benchmark task set spanning four levels of complexity and conducted experiments in a real Misty robot environment. Extensive evaluations demonstrate that AutoMisty not only consistently generates high-quality code but also enables precise code control, significantly outperforming direct reasoning with ChatGPT-4o and ChatGPT-o1. All code, optimized APIs, and experimental videos will be publicly released through the webpage: AutoMisty.
@@ I INTRODUCTION
Social robots are becoming increasingly prevalent in homes, workplaces, and educational settings, providing companionship, assistance, and engagement across various applications. Misty, a versatile social robot, is gaining adoption for interactive and assistive roles. For example, it can serve as a mental well-being coach [1], a caregiver for the elderly [2], or a learning companion for children [3]. As these robots evolve to support diverse interactions, customizing their behavior to align with individual preferences remains a challenge. Despite the availability of open APIs and developer-friendly frameworks, personalization often requires advanced programming skills, making it inaccessible to non-technical users [4, 5].
Large Language Models (LLMs) are emerging as a promising solution for enhancing human-robot interaction due to their advanced code generation capabilities. Recent research has explored leveraging LLMs to improve programming systems, reducing the need for extensive coding skills [6, 7]. However, a fully automated, end-to-end system has yet to be realized. At the same time, existing LLM-driven frameworks face significant challenges, including reasoning limitations, hallucinations, and execution reliability issues, leading to frequent errors in robot programming and action execution [8]. Additionally, the absence of robust feedback mechanisms and API integration often results in unreliable task execution, as many LLMs struggle to retain essential contextual information over prolonged interactions [9]. Recent advancements in multi-agent LLM frameworks have attempted to address these issues by incorporating structured reasoning and collaborative planning for more reliable task execution [10]. However, further optimization is still required to ensure adaptability and responsiveness within a fully end-to-end system.
@@
To tackle these challenges, we introduce AutoMisty, illustrated in Fig.1, a novel multi-agent LLM framework that enhances human-robot interaction. It enables autonomous task planning, task assignment, self-reflection, and executable code generation, all while adapting to user preferences. Our key contributions are as follows:
  • We present AutoMisty, the first multi-agent system designed for natural language-driven robot programming. It autonomously formulates task plans, performs reasoning, and generates executable code, allowing Misty to execute complex instructions with user-specific adaptability.
  • We design a two-layer reasoning mechanism for better decision-making and execution reliability: (i) self-reflection for iterative improvement and (ii) human-in-the-loop feedback for continuous adaptation. Additionally, we optimize the APIs to enhance interpretability for LLMs.
@@ II RELATED WORKS
Advancements in LLMs and natural language-driven programming have accelerated interactive multi-agent systems for robotics. AutoMisty leverages AI-driven code generation, multi-agent collaboration, and interactive programming to decompose human instructions into executable robotic control.
@@
Framework for Multi-Agent LLM Systems: The Multi-Agent LLM Framework abstracts LLM models into distinct agents, each assigned specialized roles, making it easier to design systems with multiple collaborative agents. Existing frameworks facilitate multi-agent interactions [34, 35], simulate software development workflows [36, 37], or generate autonomous agents [38]. Automisty is built upon the Autogen framework as its foundational implementation [34].
@@ III PROBLEM FORMULATION
Given a natural language task $\mathcal{I}$, AutoMisty aims to generate executable MistyCode code $C$ to accomplish the given task. Initially, the system performs semantic analysis and requirement delineation on $\mathcal{I}$, then breaks it down into several subtasks $T=T_{1},T_{2},...,T_{k}$. If a subtask is too complex, it can be further subdivided until each subtask is sufficiently atomic to execute. The system then leverages a set of $N$ large language model Agents, denoted by $A$, to handle each subtask. Each Agent in this set specializes in a different skill set to handle a specific subtask. A unique assignment function then maps each subtask $T_{j}$ to an Agent that meets its skill requirements. This is defined as:
$\phi:\{\,T_{1},\dots,T_{k}\}\;\to\;A$
If the output of Agent $A_{j}$ is required as input for Agent $A_{i}$, a directed edge $(A_{j}\to A_{i})$ is added to a directed acyclic graph $G$, which represents this multi-agent collaboration. As $G$ is acyclic, a topological sort produces the execution sequence:
@@
ensuring that for any Agent in the sequence, all of its upstream dependencies have finished before it executes. For the $k$-th Agent in the execution order, $A_{i_{k}}$, its input condition, $\Delta(A_{i_{k}})$ comprises all the code or contextual information $C$ generated by its upstream Agents. We then define the code generation function
$\Gamma(A_{i_{k}}\,\mid\,\Delta(A_{i_{k}}))\;=\;C_{i_{k}},$
where $C_{i_{k}}$ is the code that $A_{i_{k}}$ produces upon completing its subtask. Once the last agent in the sequence is reached, its output becomes the final MistyCode $C$ that completes the task $\mathcal{I}$.
@@ IV METHODOLOGY
This section outlines the AutoMisty methodology. We first introduce the optimized Misty APIs and their assignments, which are essential for enhancing LLM comprehension and helping agents focus on their designated tasks. Next, we present the two-layer optimization mechanism within each agent, ensuring reliable and well-structured code. Finally, we describe the collaboration among the four agents utilized in our proposed method and detail their specific functions. Fig. 2 illustrates this approach.
@@ IV-A Optimized Misty APIs
We observe several issues with Misty’s APIs, including unclear function descriptions, missing parameter definitions, and incomplete interfaces, which often lead to errors and hallucinations during execution. To ensure high-quality and efficient code generation, an optimized API is essential.
Inspired by Song et al. [39], we find that detailed code annotations greatly enhance LLM accuracy in code generation and execution. We first analyze Misty’s capabilities, identifying three core task categories: action, touch, and audiovisual processing. We then refine and restructure the APIs, improving audio-video data flow and integrating OpenAI’s Whisper model for real-time translation, enabling Misty to “see” and “hear.” Additionally, we encapsulate LLM logic, equipping Misty with basic reasoning and understanding. This process resulted in 136 optimized APIs, for which we provide comprehensive documentation on input/output formats, parameter semantics, application scenarios, and complex functionalities, helping to reduce errors and enhance reliability. Finally, these APIs are assigned to specific subtask agents to improve focus and reduce the likelihood of hallucinations.
@@ IV-B Two-layer Optimization Mechanism
AutoMisty employs a Two-Layer Optimization Mechanism within each agent to enhance reasoning, decision-making, and task execution. When an agent is triggered, it activates a Drafter, which listens to Layer 1’s dialog. Layer 1 utilizes a Self-Reflective Feedback Mechanism, where the Designer iteratively refines the draft using In-Context Learning (ICL), while the Critic evaluates its accuracy and feasibility. If revisions are needed, feedback loops between these two agents until the Critic approves the optimal solution. Once this phase is finalized, the Drafter initiates Layer 2, a human-in-the-loop process where the user provides direct feedback on the current results. If any aspects are deemed unsatisfactory, the Drafter reactivates Layer 1, iteratively refining the output until the user grants final approval. Once the user is satisfied, the agent concludes the current subtask.
The optimization mechanism includes a memory module that enhances the agent’s teachability, allowing it to learn from past successes. When a user saves a successful task, the system refines future code generation while preserving user preferences, improving efficiency and consistency.
@@ IV-C Multi-Agent Collaboration Framework
The Planner Agent serves as the central coordinator, responsible for analyzing high-level instructions, breaking down tasks into manageable subtasks, and assigning them to the appropriate agents. It generates an action plan with an optimized execution order and ensures coherence before proceeding. It also seeks human validation to align execution with user intent. Once the final plan is generated, the Planner routes it to subtask agents in the designated execution order. The Action, Touch, and Audiovisual Agents execute specialized tasks within their respective domains. Each agent engages in fine-grained reasoning, analyzing relevant APIs, and generating executable code, optionally referencing prior stored memory.
In addition, each agent is designed to isolate their reasoning from each other. This approach allows agents to operate collectively yet independently, ensuring that each agent focuses only on its designated subtask while relying on structured interactions to exchange relevant information. This isolation prevents any agent from being overwhelmed with excessive context, thus improving efficiency and decision-making accuracy.
IV-D System Verification and User Proxy
Once the MistyCode has been generated, AutoMisty verifies the code in a local environment before forwarding it to Misty for execution. If compilation or runtime errors occur, the system captures exception details and feeds them back to the Designer, enabling debugging and refinement within an updated context that includes error messages and partial execution results. This execution-feedback loop repeats until the code runs successfully in the local environment. Once the code passes verification, Misty executes the corresponding instructions. Inspired by AutoGen [34], we utilize the User Proxy to proxy user input to the system, allowing the user to observe the robot’s actual behavior and propose further refinements as needed.
@@ V-A Benchmark Dataset
To evaluate AutoMisty’s performance and compare it with baseline methods, we created a benchmark dataset for natural language-based task code generation and execution with the Misty robot. We invited 10 people (3 engineers and 7 non-engineers) to suggest tasks for Misty, merging overlapping ones. The dataset includes tasks of varying complexity, grouped into four categories:
  • Elementary Tasks include basic commands such as raising the left or right hand, adjusting head angles, and controlling movement speed. The primary goal is to test the user’s ability to precisely control the robot through conversation.
  • Simple Tasks are tasks that can be completed using a single agent with minimal creativity. The solution is straightforward, and execution is direct.
  • Compound Tasks require multiple agents to collaborate in completing the task. The instructions are clear, and the interaction between agents follows a well-defined structure.
  • Complex Tasks are highly abstract and vague instructions that require multiple agents working in tandem. The system must autonomously design LLM prompts to accomplish the task. Solving the task demands significant creativity, potentially involving external tool integration or even the design of new tools.
The dataset consists of 28 tasks in total: 9 elementary tasks, 7 simple tasks, 4 compound tasks, and 8 complex tasks. Table I provides a list of tasks under each category.
@@ V-B Evaluation Metrics
  • US measures user satisfaction with the generated code. If an interaction’s instructions are misinterpreted or not followed, a penalty of -1 is applied. The maximum possible score is 10.
@@ VI RESULTS AND DISCUSSION
Teachability Assessment: Teachability refers to an agent’s ability to learn from past successes. To evaluate this capability in our system, we designed an experiment based on user-specific emotional preferences. Since users perceive emotions uniquely, we stored four well-recognized emotion types in the memory. The user then inputs synonymous emotions, and we assess whether the system correctly retrieves the corresponding anchor emotion.Table V summarizes the results. Although AutoMisty misclassifies some emotions, it accurately retrieves well-defined ones like ”Rage” and ”Grief,” indicating its teachability.
@@ VII CONCLUSION
AutoMisty is the first system to break the technical bottleneck, enabling non-programmers to generate executable code directly through conversational instructions. Our design significantly enhances code control, accuracy, and consistency, achieving a 100% task success rate across all complexity levels, whereas ChatGPT-o1 achieves only 75% on complex tasks. While we cover 28 tasks, our system demonstrates strong generalization capabilities for broader open-domain tasks. By leveraging context optimization strategies, such as API semantic anchoring, the entire framework operates solely through in-context learning, enabling low-cost and efficient migration to other API-driven social robotics. In the future, we will expand our work to include collaboration and interaction between multiple social robots and humans.

### v2

@@ Abstract
The social robot’s open API allows users to customize open-domain interactions. However, it remains inaccessible to those without programming experience. We introduce AutoMisty, the first LLM-powered multi-agent framework that converts natural-language commands into executable Misty robot code by decomposing high-level instructions, generating sub-task code, and integrating everything into a deployable program. Each agent employs a two-layer optimization mechanism: first, a self-reflective loop that instantly validates and automatically executes the generated code, regenerating whenever errors emerge; second, human review for refinement and final approval, ensuring alignment with user preferences and preventing error propagation. To evaluate AutoMisty’s effectiveness, we designed a benchmark task set spanning four levels of complexity and conducted experiments in a real Misty robot environment. Extensive evaluations demonstrate that AutoMisty not only consistently generates high-quality code but also enables precise code control, significantly outperforming direct reasoning with ChatGPT-4o and ChatGPT-o1. All code, optimized APIs, and experimental videos will be publicly released through the webpage: AutoMisty.
@@ I INTRODUCTION
Social robots are becoming increasingly prevalent in homes, workplaces, and educational settings, providing companionship, assistance, and engagement across various application [1]. Misty, a versatile social robot, is gaining adoption for interactive and assistive roles. For example, it can serve as a mental well-being coach [2], a caregiver for the elderly [3], or a learning companion for children [4]. As these robots evolve to support diverse interactions, customizing their behavior to align with individual preferences remains a challenge. Despite the availability of open APIs and developer-friendly frameworks, personalization often requires advanced programming skills, making it inaccessible to non-technical users [5, 6].
Large Language Models (LLMs) are emerging as a promising solution for enhancing human-robot interaction due to their advanced code generation capabilities. Recent research has explored leveraging LLMs to improve programming systems, reducing the need for extensive coding skills [7, 8]. However, a fully automated, end-to-end system has yet to be realized. At the same time, existing LLM-driven frameworks face significant challenges, including reasoning limitations, hallucinations, and execution reliability issues, leading to frequent errors in robot programming and action execution [9]. Additionally, the absence of robust feedback mechanisms often leads to undesired task behavior, resulting in error propagation. This issue is further exacerbated by the context length limitations of many large language models, which struggle to retain essential information throughout prolonged interactions [10]. Recent work in robotics has explored multi-agent collaboration systems based on large language models (LLMs), incorporating structured reasoning, task planning, and dedicated agent collaboration to improve task execution reliability [11]. However, further refinement is needed to ensure adaptability and responsiveness in fully end-to-end systems.
@@
To tackle these challenges, we introduce AutoMisty (illustrated in Fig.1), a novel multi-agent LLM framework designed to generate reliable Misty robot control code from natural language task descriptions. The framework supports abstract task understanding and planning, subtask assignment, executable code generation, and real-time code deployment to the Misty robot. Furthermore, users can provide feedback based on the robot’s actual performance, enabling human-in-the-loop collaboration to improve code quality and execution outcomes. Our key contributions are as follows:
  • We present AutoMisty, the first multi-agent system designed for natural language-driven Misty robot programming. It autonomously formulates task plans, performs reasoning, and generates executable code, allowing Misty to execute complex instructions with user-specific adaptability.
  • We design a two-layer mechanism to ensure reliable and user-aligned task execution: (i) a self-reflective loop that validates and executes generated code in real time, automatically regenerating it when errors are detected; and (ii) a human feedback loop for iterative refinement and final approval, ensuring alignment with user intent and preventing error propagation. In addition, we optimize the Misty control APIs to improve interpretability and call reliability for LLMs.
@@ III PROBLEM FORMULATION
Given a natural language task $\mathcal{I}$, AutoMisty aims to generate executable MistyCode code $C$ to accomplish the given task. Initially, the system performs semantic analysis and requirement delineation on $\mathcal{I}$, then breaks it down into several sub-tasks $T=T_{1},T_{2},...,T_{k}$. If a sub-task is too complex, it can be further subdivided until each sub-task is sufficiently atomic to execute. The system then leverages a set of $N$ LLM agents, denoted as $A=\{A_{1},A_{2},\ldots,A_{N}\}$, where each $A_{i}$ specializes in handling a specific type of sub-task. An assignment function then maps each sub-task to its corresponding agent, whose capabilities match the subtask’s requirements. This is formally defined as:
$\phi:\{\,T_{1},\dots,T_{k}\}\;\to\;\{A_{1},A_{2},\dots,A_{N}\}$
The overall task can be naturally formulated as a directed acyclic graph (DAG), where agents are nodes and dependency relations define edges. If the output of agent $A_{j}$ is required as input for agent $A_{i}$, a directed edge $(A_{j}\to A_{i})$ is added to a directed acyclic graph $G$, which represents this multi-agent collaboration. As $G$ is acyclic, a topological sort produces the execution sequence:
@@
ensuring that for any agent in the sequence, all of its upstream dependencies have finished before it executes. For the $k$-th agent in the execution order, $A_{i_{k}}$, its input includes the upstream-generated code $\Delta(A_{i_{k}})$ and its own LLM context $\text{Ctx}(A_{i_{k}})$, such as system prompts and subtask descriptions. The code generation function is defined as:
$\Gamma(A_{i_{k}}\,\mid\,\text{Ctx}(A_{i_{k}}),\,\Delta(A_{i_{k}}))\;=\;C_{i_{k}},$
where $C_{i_{k}}$ is the code generated by $A_{i_{k}}$ for its assigned subtask. The final Misty Code output $C$ corresponds to the output of the last agent in the topological execution sequence, completing the original task $\mathcal{I}$.
@@ IV METHODOLOGY
This section outlines the AutoMisty methodology. We first present the optimized Misty API design and its functional partitioning across agents. This design enhances LLM comprehension, mitigates context window issues that prevent accurate API selection and internal code modification, and helps agents focus on their assigned subtasks. We then introduce a unified two-layer architecture shared by all agents to ensure reliable and well-structured code generation. Finally, we describe the collaboration among the four core agents and their respective roles.Fig. 2 illustrates this approach.
@@ IV-A Optimized Misty APIs
We identified several issues in Misty’s original APIs, such as vague descriptions, missing parameters, and incomplete interfaces, often causing execution errors or hallucinations. To enable high-quality code generation, we optimized the API set. Following insights from Song et al. [35], we found that detailed annotations significantly improve LLM performance. We analyzed Misty’s core capabilities—action, touch, and audiovisual processing—and restructured the APIs accordingly. This includes refining audio-video flow, integrating Whisper for real-time translation, and enabling basic verbal reasoning via LLM/VLM. The final set includes 136 optimized APIs with clear documentation on I/O formats, parameters, use cases, and complex functions. These APIs are embedded in system prompts and routed to agents for task-specific code generation.
@@ IV-B Two-layer Optimization Mechanism
AutoMisty integrates a Two-Layer Optimization Mechanism within each agent to enhance reasoning, decision-making, and task execution. Upon activation, the agent triggers a Drafter that manages Layer 1, which adopts a Self-Reflective Feedback Mechanism: the Designer generates initial code via In-Context Learning (ICL) and iteratively refines it based on feedback from the Critic and, when available, the user. The Critic evaluates the code’s correctness and feasibility. If the solution is unsatisfactory, the Designer and Critic engage in a feedback loop until the Critic approves the optimized version. These interactions are orchestrated through sequential prompting, enabling structured, stepwise refinement within the LLM context.
The Designer also incorporates a memory module based on retrieval-augmented generation (RAG). When users save a successful task, the system stores the verified code and preferences. For future similar tasks, the module retrieves relevant examples before generation, providing few-shot demonstrations that align with user intent and prior successes.
After Layer 1, the Drafter proceeds to Layer 2—the human-in-the-loop phase. The UserProxy component integrates a compiler tailored to the Misty runtime environment and automatically executes the generated code. If execution errors occur, they are appended to the context and returned to the Designer for revision. Once the code runs successfully and is deployed on the Misty robot, the user can provide feedback on its behavior via typed input. If the user is not satisfied, the Drafter reactivates Layer 1 and incorporates the human feedback into the context to further refine the output. The subtask is only considered complete after the user explicitly confirms success.
@@ IV-C Multi-agent Collaboration Framework
The Planner agent serves as the central coordinator, interpreting vague high-level human instructions, enriching them with necessary details, decomposing tasks into manageable subtasks, and assigning them to the appropriate agents.It generates a coherent, optimized execution plan, validates it with the user, and routes subtasks to agents accordingly.
Action, Touch, and Audiovisual agents each operate within their respective functional domains. The Action agent controls Misty’s speech output, gestures, and head movements. The Touch agent processes physical interactions from users and automatically translates them into system-recognizable responses. The Audiovisual agent handles visual and auditory inputs, integrating both a large language model (LLM) and a vision-language model (VLM), enabling Misty to reason and respond based on what it sees and hears, thereby providing fundamental multimodal understanding and interaction capabilities.
In addition, each agent is designed to access its own internal context and is unaware of the intermediate reasoning processes of other agents. At most, it receives the final, verified code output from the preceding agent, which must be embedded verbatim into its own generated code. This design prevents context overflow caused by long cross-agent context histories and ensures that each agent remains focused solely on its assigned subtask, thereby improving both efficiency and system robustness.
@@ V-A Benchmark Dataset
To evaluate AutoMisty’s performance, we constructed a natural language task suite for the Misty robot. Tasks are freely described without constraints on format or detail, allowing highly abstract instructions (e.g., “do a rock dance”). We collected task proposals from 10 participants (3 engineers, 7 non-engineers), merged overlapping entries, and asked each participant to define their own success criteria. The final suite covers tasks of varying complexity, grouped into four categories.
  • Elementary Tasks are basic low-level actions executable via a single API call, such as raising a hand, adjusting head orientation, or setting speed. They evaluate the system’s ability to map natural language commands to atomic robot controls.
  • Simple Tasks are tasks that can be completed by a single agent, with clear solutions and straightforward execution.
  • Compound Tasks require collaboration among multiple agents. The instructions are clear and Misty can complete the task independently without relying on an LLM as the central reasoning component.
  • Complex Tasks involve abstract and ambiguous instructions that require coordination among multiple agents and rich human-robot interaction. The system must autonomously design the prompts for the LLM or VLM acting as Misty’s cognitive core. Solving these tasks demands strong abstraction understanding and planning capabilities.
The task suite consists of 28 tasks in total: 9 elementary tasks, 7 simple tasks, 4 compound tasks, and 8 complex tasks. Table I provides a list of tasks under each category.
@@ V-B Evaluation Metrics
  • US measures user satisfaction with the generated code. Users were asked to document their interactions while using AutoMisty. If the system misinterpreted or failed to follow an instruction, a penalty of 1 point was applied. The maximum score for this metric is 10.
@@ VI RESULTS AND DISCUSSION
Teachability Assessment: Teachability refers to an agent’s ability to learn from prior successes and adapt to user preferences. AutoMisty supports this by using a vector database for retrieval, where each key is a compressed subtask description ((5–10 word high-level summary generated via an LLM)), and the value is user-verified code. These examples reflect both prior successes and user preferences, guiding future generations toward more personalized outputs.
We evaluated this capability through an emotion-grounded experiment. We first injected 10 irrelevant tasks into the memory and stored four anchor emotions (e.g., Rage, Grief), each mapped to expressive Misty behaviors involving motion, facial expressions, and sound. Users then input synonymous or vague emotion terms. For each query, we retrieved only the top-1 result from memory. As shown in Table V, AutoMisty correctly retrieved the corresponding emotion class code for clearly defined inputs, demonstrating teachability. Additionally, the generated code for each emotion exhibited consistent behavioral patterns.
@@ VII CONCLUSION
AutoMisty is the first system to break the technical bottleneck, enabling non-programmers to generate executable code directly through conversational instructions. Our design significantly improves code controllability, accuracy, and consistency, demonstrating strong robustness across tasks of varying complexity. Although the current task suite includes only 28 examples, the system exhibits strong generalization capabilities for broader open-domain tasks. The entire AutoMisty framework is built on in-context learning, enabling low-cost and efficient transfer to other API-driven social robotics platforms.
Acknowledgment
This material is based upon work supported under the AI Research Institutes program by the U.S. National Science Foundation and the Institute of Education Sciences, U.S. Department of Education, through Award # 2229873—National AI Institute for Exceptional Education. Any opinions, findings and conclusions, or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of the National Science Foundation, the Institute of Education Sciences, or the U.S. Department of Education.